# Confucius3-Math: A Lightweight High-Performance Reasoning LLM for Chinese K-12 Mathematics Learning

### v1

@@ 1 Introduction
Reasoning capabilities have been the focus of the latest Large Language Model (LLM) developments. OpenAI’s o1 model (OpenAI, 2024) represents a major breakthrough in advancing LLM’s abilities in solving complex, multi-step tasks such as math and coding, igniting mainstream awareness of reasoning-specific LLMs. o1 is known to acquire its performance via Test Tine Compute (TTC). However, neither the model weights nor the technical details of o1, or its sequels OpenAI (2025), are released. TTC is a well-known technique to boost model’s performance by allowing the model to “think” or compute more at inference time and has been applied successfully in many reasoning tasks, mostly in games such as Backgammon (Tesauro, 1994), poker (Brown & Sandholm, 2017), and Go (Silver et al., 2016) etc. It takes on many forms from simple consensus/majority vote to beam search and Monte-Carlo Tree Search (MCTS) (Metropolis & Ulam, 1949), depending on the application. Jones (2021) studies scaling laws between train time and test time, clarifying the benefit of TTC, also in the game setting.
@@
    Confucius3-Math demonstrates the feasibility of building strong and practical reasoning models in a particular domain at low cost. Specifically, (1) Confucius3-Math achieves SOTA performance on a number of K-12 math benchmarks; (2) The training of Confucius3-Math costs only \$26K; and (3) its inference performance is about 15$\times$ that of DeepSeek R1.
@@
    We introduce Targeted Entropy Regularization, Recent Sample Recovery and Policy-Specific Hardness Weighting. These innovations encompass a new entropy regularization, a novel data scheduling policy, and an improved group-relative advantage estimator. They significantly stabilize the RL training, improve data efficiency, and boost performance. Recent Sample Recovery and Policy-Specific Hardness Weighting, one novel data scheduling policy and one improved group-relative advantage estimator, respectively, that significantly improves data efficiency, stabilizes the RL training process, and boost performance.
@@ 2 Approach
For post-training, we contemplate a number of choices, including RL, SFT, distillation, and their combinations. For enhancing small models, past works have shown their effectiveness in different situations. For example, DeepSeek R1 (DeepSeek-AI et al., 2025) shows that distillation obtains much larger performance gains than post-training the same model via RL. The latest Qwen3 series also use strong-to-weak distillation to optimize their lightweight models (Qwen3-0.6B, 1.7B, 4B, 8B, and 14B) and they show that a combination of off-policy/on-policy distillation achieves significantly better performance than RL (Yang et al., 2025).
@@ 4 Base Model Selection
We discover that all display certain level of trainability, with DeepSeek-R1-Distill-Qwen-14B performing the best. And indeed, the model’s outcome is correlated with its initial entropy. We postulate that the policy entropy could be a critical metric when determining whether a model can serve as an initial model for RL optimization. Interestingly, among the models, Confucius-o1-14B, which is trained from Qwen2.5-14B-Instruct, shows a lower entropy than its base, which appears to be caused by its training method: DPO without entropy regularization.
@@ 6.1 Targeted Entropy Regularization
However, from our experience, entropy seems to be a delicate matter. LLMs are complex machines with output distribution over a large space of multilingual tokens. The maximum entropy principle appears to have double effects. One major issue we encounter during the training process is the “mixed language problem”, a phenomenon that also arises with DeepSeek R1 (DeepSeek-AI et al., 2025). This phenomenon appears to be correlated with high entropy, which may encourage the model to substitute tokens with the same meaning from different languages. We tried adding linguistic constraints to the reward, but it was ineffective.
@@ 7 Evaluation
Results. The main results are presented in table 3. It shows the accuracy rates of the models on a number of benchmarks. The same data is also shown visually in figure 1 in the beginning of this paper. The first 4 rows in table 3, as well as the left part of figure 1 with pink background, are the benchmarks pertinent to general K-12 learning, while the rest are competition-level benchmarks. Confucius3-Math outperforms all other models on all but two benchmarks, on which Confucius3-Math’s performance is comparable to the strongest models. In particular, on the CK12-MATH benchmark, which is the most important focus of our application scenario, Confucius3-Math leads DeepSeek R1 by 3.5 points. Qwen3-14B, as a 14B model, also performs quite well on some tasks. However, Qwen3-14B starts with a much stronger initial model, consumes more data, and relies on more powerful teacher models. We believe Confucius3-Math’s RL solution points to an alternative, and maybe more effective, technical route. As mentioned earlier, we still have a large amount of unused data relevant to this scenario and continued improvement is almost certain. This high accuracy, plus its low inference cost (see section 8), makes Confucius3-Math a viable LLM solution in our actual business operations.
@@ 8 Inference Performance
The Confucius3-Math model outperforms DeepSeek R1 671B on K-12 math reasoning tasks. With only 14B parameters, it can be deployed with much lower cost, a crucial factor affecting the sustainability of a service. We conduct tests to verify Confucius3-Math’s inference efficiency under two configurations: (1) general throughput measurements with abundant resources, and (2) low-resource deployment feasibility test. For (1), we run both R1 and Confucius3-Math on the same high-performance hardware and measure their throughput, in terms of tokens per second. Specifically, we use a machine with the following hardware components:
@@
The models are deployed using the vLLM 0.9.0 framework, and CUDA (version 12.8). Both prefix caching and chunk filling are enabled during these tests. We choose FP8 as the model quantization level. For DeepSeek R1, a parallel strategy with TP = 8 is employed. Confucius3-Math inference is run on a single GPU while DeepSeek R1 671B on 8, to make sure R1 is deployed with adequate resources. We scale the results of Confucius3-Math by a factor of 8 for the two to be comparable. All other parameters are configured to the default settings of the vLLM framework.
A model with low resource requirement is advantageous, as it has more application opportunities. For this test we deploy Confucius3-Math with FP8 precision on a single GeForce RTX 4090D GPU with 24GB memory. Again we scale the numbers by 8 for comparability. Note that it is not possible to run DeepSeek R1 in this setting at all.
@@
Results are shown in figure 5. For each run, solid line represents total throughput, in tokens per second, while dashed line is output throughput, also in tokens per second. DeepSeek R1 and Confucius3-Math’s performances on H800 are shown as green dots and red triangles, respectively. Confucius3-Math’s throughputs on 4090D are represented as red $\times$.
A few patterns are obvious. Firstly, Confucius3-Math outperforms R1 by large margins at all request rates. Secondly, R1 saturates quickly at 1,513 and 1,631 tokens per second on output and total throughput, respectively, at QPS = 1. This means that the system reaches its processing capacity at very low workload. Confucius3-Math, on the other hand, starts with 13,182 and 14,211 tokens per second at QPS = 1, scales more gracefully as QPS grows. It reaches its peak throughput at QPS 4, with a total throughput of 31,994. Compared with DeepSeek R1, this is a 15.8$\times$ speedup.
Thirdly, even running on low-resource hardware, Confucius3-Math surpasses DeepSeek R1’s performance on H800. The model achieves impressive inference input and output throughputs of up to 4,596 and 4,956 tokens per second, respectively, more than double those of R1.
@@ 9 Feasibility and Cost
An alternative solution, strong-to-weak distillation, is a commonly used method to enhance small models. It is actually more complex than it appears and may involve hidden costs. Acquiring outputs from teacher models can be done via API for proprietary models or hosting inference services for open models. Neither is free. And the more powerful the teacher, the more expensive. In particular, on-policy distillation, such as that used in training Qwen3-14B (Yang et al., 2025), requires more frequent and real-time access to the teacher model. This is only possible when one has access to the model weights, further narrowing down one’s choices. Again, powerful models are usually large and hosting them is not a negligible cost. For example, on-policy distillation using DeepSeek R1 as the teacher model needs, in addition to training GPUs, at least one extra H800-like server (to host the teacher) to run, let alone efficiently.

### v2

@@ 1 Introduction
Reasoning capabilities have been the focus of the latest Large Language Model (LLM) developments. OpenAI’s o1 model (OpenAI, 2024) represents a major breakthrough in advancing LLM’s abilities in solving complex, multi-step tasks such as math and coding, igniting mainstream awareness of reasoning-specific LLMs. o1 is known to acquire its performance via Test Time Compute (TTC). However, neither the model weights nor the technical details of o1, or its sequels OpenAI (2025), are released. TTC is a well-known technique to boost model’s performance by allowing the model to “think” or compute more at inference time and has been applied successfully in many reasoning tasks, mostly in games such as Backgammon (Tesauro, 1994), poker (Brown & Sandholm, 2017), and Go (Silver et al., 2016) etc. It takes on many forms from simple consensus/majority vote to beam search and Monte-Carlo Tree Search (MCTS) (Metropolis & Ulam, 1949), depending on the application. Jones (2021) studies scaling laws between train time and test time, clarifying the benefit of TTC, also in the game setting.
@@
    Confucius3-Math demonstrates the feasibility of building strong and practical reasoning models in a particular domain at low cost. Specifically, (1) Confucius3-Math achieves SOTA performance on a number of K-12 math benchmarks; (2) The training of Confucius3-Math costs only \$26K; and (3) its inference performance is about 15$\times$ that of DeepSeek-R1.
@@
    We introduce Targeted Entropy Regularization, Recent Sample Recovery and Policy-Specific Hardness Weighting. These innovations encompass a new entropy regularization, a novel data scheduling policy, and an improved group-relative advantage estimator. They significantly stabilize the RL training, improve data efficiency, and boost performance.
@@ 2 Approach
For post-training, we contemplate a number of choices, including RL, SFT, distillation, and their combinations. For enhancing small models, past works have shown their effectiveness in different situations. For example, DeepSeek-R1 (DeepSeek-AI et al., 2025) shows that distillation obtains much larger performance gains than post-training the same model via RL. The latest Qwen3 series also use strong-to-weak distillation to optimize their lightweight models (Qwen3-0.6B, 1.7B, 4B, 8B, and 14B) and they show that a combination of off-policy/on-policy distillation achieves significantly better performance than RL (Yang et al., 2025).
@@ 4 Base Model Selection
We discover that all display certain level of trainability, with DeepSeek-R1-Distill-Qwen-14B performing the best. And indeed, the model’s outcome is correlated with its initial entropy. We postulate that the policy entropy could be an importance metric when determining whether a model can serve as an initial model for RL optimization. Interestingly, among the models, Confucius-o1-14B, which is trained from Qwen2.5-14B-Instruct, shows a lower entropy than its base, which appears to be caused by its training method: DPO without entropy regularization.
@@ 6.1 Targeted Entropy Regularization
However, from our experience, entropy seems to be a delicate matter. LLMs are complex machines with output distribution over a large space of multilingual tokens. The maximum entropy principle appears to have double effects. One major issue we encounter during the training process is the “mixed language problem”, a phenomenon that also arises with DeepSeek-R1 (DeepSeek-AI et al., 2025). This phenomenon appears to be correlated with high entropy, which may encourage the model to substitute tokens with the same meaning from different languages. We tried adding linguistic constraints to the reward, but it was ineffective.
@@ 7 Evaluation
Results. The main results are presented in table 3. It shows the accuracy rates of the models on a number of benchmarks. The same data is also shown visually in figure 1 in the beginning of this paper. The first 4 rows in table 3, as well as the left part of figure 1 with pink background, are the benchmarks pertinent to general K-12 learning, while the rest are competition-level benchmarks. Confucius3-Math outperforms all other models on all but two benchmarks, on which Confucius3-Math’s performance is comparable to the strongest models. In particular, on the CK12-MATH benchmark, which is the most important focus of our application scenario, Confucius3-Math leads DeepSeek-R1 by 3.5 points. Qwen3-14B, as a 14B model, also performs quite well on some tasks. However, Qwen3-14B starts with a much stronger initial model, consumes more data, and relies on more powerful teacher models. We believe Confucius3-Math’s RL solution points to an alternative, and maybe more effective, technical route. As mentioned earlier, we still have a large amount of unused data relevant to this scenario and continued improvement is almost certain. This high accuracy, plus its low inference cost (see section 8), makes Confucius3-Math a viable LLM solution in our actual business operations.
@@ 8 Inference Performance
The Confucius3-Math model outperforms DeepSeek-R1 671B on K-12 math reasoning tasks. With only 14B parameters, it can be deployed with much lower cost, a crucial factor affecting the sustainability of a service. We conduct tests to verify Confucius3-Math’s inference efficiency under two configurations: (1) general throughput measurements with abundant resources, and (2) low-resource deployment feasibility test. For (1), we run both R1 and Confucius3-Math on the same high-performance hardware and measure their throughput, in terms of tokens per second. Specifically, we use a machine with the following hardware components:
@@
The models are deployed using the vLLM 0.9.0 framework, and CUDA (version 12.8). Both prefix caching and chunk filling are enabled during these tests. We choose FP8 as the model quantization level. For DeepSeek-R1, a parallel strategy with TP = 8 is employed. Confucius3-Math inference is run on a single GPU while DeepSeek-R1 671B on 8, to make sure R1 is deployed with adequate resources. We scale the results of Confucius3-Math by a factor of 8 for the two to be comparable. All other parameters are configured to the default settings of the vLLM framework.
A model with low resource requirement is advantageous, as it has more application opportunities. For this test we deploy Confucius3-Math with FP8 precision on a single GeForce RTX 4090D GPU with 24GB memory. Again we scale the numbers by 8 for comparability. Note that it is not possible to run DeepSeek-R1 in this setting at all.
@@
Results are shown in figure 5. For each run, solid line represents total throughput, in tokens per second, while dashed line is output throughput, also in tokens per second. DeepSeek-R1 and Confucius3-Math’s performances on H800 are shown as green dots and red triangles, respectively. Confucius3-Math’s throughputs on 4090D are represented as red $\times$.
A few patterns are obvious. Firstly, Confucius3-Math outperforms R1 by large margins at all request rates. Secondly, R1 saturates quickly at 1,513 and 1,631 tokens per second on output and total throughput, respectively, at QPS = 1. This means that the system reaches its processing capacity at very low workload. Confucius3-Math, on the other hand, starts with 13,182 and 14,211 tokens per second at QPS = 1, scales more gracefully as QPS grows. It reaches its peak throughput at QPS 4, with a total throughput of 31,994. Compared with DeepSeek-R1, this is a 15.8$\times$ speedup.
Thirdly, even running on low-resource hardware, Confucius3-Math surpasses DeepSeek-R1’s performance on H800. The model achieves impressive inference input and output throughputs of up to 4,596 and 4,956 tokens per second, respectively, more than double those of R1.
@@ 9 Feasibility and Cost
An alternative solution, strong-to-weak distillation, is a commonly used method to enhance small models. It is actually more complex than it appears and may involve hidden costs. Acquiring outputs from teacher models can be done via API for proprietary models or hosting inference services for open models. Neither is free. And the more powerful the teacher, the more expensive. In particular, on-policy distillation, such as that used in training Qwen3-14B (Yang et al., 2025), requires more frequent and real-time access to the teacher model. This is only possible when one has access to the model weights, further narrowing down one’s choices. Again, powerful models are usually large and hosting them is not a negligible cost. For example, on-policy distillation using DeepSeek-R1 as the teacher model needs, in addition to training GPUs, at least one extra H800-like server (to host the teacher) to run, let alone efficiently.